# Data Trade and Consumer Privacy

### v2

@@ 2 Model
The intermediary plays in both service and data markets by offering a contract of data collection and data trade. In the service market the intermediary is a monopolist who provides consumers service to collect consumers data on willingness to pay. The intermediary sells data via a data offer and the producer buys data for product pricing decision. The intermediary is the unique data source. Its profit comes from service provision and data sale.
By the revelation principle (Myerson, 1979), the intermediary designs a direct mechanism which specifies: (1) a service menu for consumers including the service provision $x:\{L,H\}\to\mathcal{X}$ where $\mathcal{X}$ is the set of service; and the service fee $f:\{L,H\}\to\mathbb{R}$; and (2) a data offer for the producer consisting of the disclosure policy captured by a pair of distributions over disclosed signals about consumers types $\pi:\left\{L,H\right\}\to\Delta\mathcal{S}$ where $\mathcal{S}$ is the set of possible signals; and the lump-sum data fee $T\in\mathbb{R}$. Consumers observe all the specifications. If a consumer reports $\theta^{\prime}$, then a signal will realize according to $\pi\left(s|\theta^{\prime}\right)$ in the product market. In sum, the intermediary offers a contract of data collection and data trade of which $\left(\mathcal{X},x,f\right)$ and ($\left(\left\{\mathcal{S},\pi\right\},T\right)$ specify the part of service market and data market respectively.
@@
Consumers privately observe type $\theta$. The intermediary posts the entire contract $\{\left(\mathcal{X},x,f\right),\left(\left\{\mathcal{S},\pi\right\},T\right)\}$ including the service menu in the service market and data offer in the data market. Then consumers and the producer make acceptance decisions. If consumers accept, they report type $\theta^{\prime}$ and then get service $x_{\theta^{\prime}}$ and pay $f_{\theta^{\prime}}$. If consumers reject, they get the outside option in the service market. If the producer accepts, after data trade, in the product market a signal $s$ about consumer’s type is realized according to $\pi$. The producer observes $s$ and then sets the product price. Consumers decide whether to purchase the product or not. If the producer rejects the data offer, then the product pricing is made under the prior information $\mu_{0}$. The timing of events is summarized in Figure 2.
@@
A key modeling assumption is the perfect correlation between consumers’ taste and their willingness to pay. That is to say, $v_{L}\neq v_{H}$ in the setup. Otherwise, the data about consumers’ tastes is not valuable to the producer. The relation between consumers’ taste parameter and their willingness to pay can be represented by $\beta_{\theta}=\text{Pr}\left(v_{H}|\theta\right)=1-\text{Pr}\left(v_{L}|\theta\right)$. The assumption reads $(\beta_{H}=1,\beta_{L}=0)$ or $(\beta_{H}=0,\beta_{L}=1)$. If we relax the assumption by $\beta_{H},\beta_{L}\in\left(0,1\right)$ with $\beta_{H}\neq\beta_{L}$, the main results don’t change.
@@ 3 Preliminaries
In the product market, a consumer $\theta$ will purchase the product if and only if $v_{\theta}\geq p$ at the product price $p$. After observing a signal $s\in\mathcal{S}$ from the traded data $\pi$, the producer’s posterior belief is $\mu_{s}\equiv\text{Pr}(\theta=H|s)$ or $\text{Pr}(v=v_{H}|s)$
@@
At the time of data trade, the producer’s expected product revenue from accepting the data trade offer ($(\{\mathcal{S},\pi\},T)$ is
$R\left(\pi\right)=\mathbb{E}_{\pi}\left[\max\left\{v_{H}\mu_{s},v_{L}\right\}\right]$ (4)
The producer’s payoff is $R\left(\pi\right)-T$ if accepting the data offer.
@@
$u\left(\theta,\theta^{\prime};\pi\right)=V-\left(x_{\theta^{\prime}}-\theta\right)^{2}-f_{\theta^{\prime}}+\mathbb{E}_{\pi\left(s|\theta^{\prime}\right)}\left[\max\left\{v_{\theta}-p_{s},0\right\}\right]$ (5)
@@ 3.1 Producer Optimality
Plugging (2) into (6) (7), the obedience constraints for the producer to follow pricing recommendations are respectively
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-(1-\frac{v_{L}}{v_{H}})\mu_{0}\pi\left(l|H\right)$$\geq$ $\frac{v_{L}}{v_{H}}-\mu_{0}$ (8)
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-(1-\frac{v_{L}}{v_{H}})\mu_{0}\pi\left(l|H\right)$$\geq$ 00 0 (9)
Now we turn to the producer’s acceptance decision over the intermediary’s data trade offer ($\left(\left\{\mathcal{S},\pi\right\},T\right)$. When the producer does not accept the intermediary’s data offer, he has no more information than the prior $\mu_{0}$ and then the revenue is $R_{0}=\max\{v_{H}\mu_{0},v_{L}\}$. The producer accepts the data offer if and only if the product revenue from accepting is weakly larger than the one from rejecting
@@
where $\mathbbm{1}$ denotes the indicator function. [^13] The special case of $\mu_{0}=v_{L}/v_{H}$ is equivalent to either the case $\mu_{0}>v_{L}/v_{H}$ or $\mu_{0}<v_{L}/v_{H}$. The data trade has two effects on the producer’s revenue. The first is the change in the probability of serving the type-L consumers. With data trade the producer sells to the type-L consumers (of mass $\left(1-\mu_{0}\right)$) with probability $\pi\left(l|L\right)$. Instead, Without data trade the type-L consumers are served with probability $\mathbbm{1}_{\mu_{0}<\frac{v_{L}}{v_{H}}}$. This is the extensive market inclusion effect of data trade.
The second effect is the change in the probability of yielding surplus to the type-H consumers. With data trade the producer charges the type-H consumers (of mass $\mu_{0}$) a low price $v_{L}$ with probability $\pi\left(l|H\right)$. Without data trade the type-H consumers are charged the low price $v_{L}$ with probability $\mathbbm{1}_{\mu_{0}<\frac{v_{L}}{v_{H}}}$. This is the intensive surplus extraction effect of data trade. [^14] It is direct that when $\mu_{0}<\frac{v_{L}}{v_{H}}$, the market inclusion effect is negative: The market serving the type-L consumers shrinks. The surplus extraction effect is positive: The producer extracts the type-H consumer surplus with a larger probability. Instead, when $\mu_{0}>\frac{v_{L}}{v_{H}}$, the market inclusion effect is positive: The market serving the type-L consumers expands. the surplus extraction effect is negative: the producer extracts the type-H consumers surplus with a smaller probability.
The triangle gray set in Figure 3 illustrates the feasible data trade gain, which depends on the prior. Take $\mu_{0}>\frac{v_{L}}{v_{H}}$ as an example. Without data trade, the revenue is $\mu_{0}v_{H}$ which is at point $A$. It can go up to point $B$ under a data trade which induces posteriors $\mu_{l}$ and $\mu_{h}$. If the traded data is of full information, the intermediary extracts the maximum data trade surplus given $\mu_{0}$, which is $\mu_{0}v_{H}+\left(1-\mu_{0}\right)v_{L}$.
@@
Figure 3: The triangle gray part is the feasible data trade gain from a data trade.
@@ 3.2 Consumers Optimality
If the type-H consumers reject the service offer, they will get nothing in the service market. In this case, the producer will treat consumers as coming from the population and make product pricing decisions based on the prior $\mu_{0}$. [^15] The intermediary cannot collect consumers data via providing service if the offer is rejected, and hence the producer cannot get extra information about consumers. That is, the product price $p_{0}$ without extra data is determined by (3) with the belief $\mu_{0}$. So the type-H consumers’ continuation value of rejecting the intermediary’s contract is
@@ 3.3 Intermediary Optimality
The intermediary maximizes the total profit (1) from both the service and data markets by specifying a contract subject to the producer’s obedience conditions (8) (9) and participation constraint (10), the consumers’ incentive constraints (12) (14) and participation constraints (13) (15), and the following two feasibility conditions of data trade
@@
We make the following short notations for simplicity.
@@
The parameter $\delta^{2}$ captures the variation in consumers’ tastes, reflecting the horizontal differentiation level in the service market, while $\Delta v$ captures the variation in consumers’ willingness to pay, measuring the vertical differentiation level in the product market.
It is ready to characterize the intermediary optimal mechanism. First, we assume a binary service space for better intuitions about the role of data trade in Section 4. Second, the section 5 extends the analysis to a general service case.
4 Binary Service
This section characterizes the intermediary optimal mechanism under the assumption of binary service $\mathcal{X}=\left\{L,H\right\}$ to focus on the cross-market role of data trade. [^16] One can treat the assumption as a delegated provision of service by the intermediary to its department of marketing, while the data trade is directly controlled by the intermediary. The marketing department’s choice of service level does not coordinate with the data trade policy. This is one of the practices: the intermediary is usually not the producer of service, but recommends service.
Lemma 2.
If $\mathcal{X}=\left\{L,H\right\}$, then $x_{H}=H$ and $x_{L}=L$.
@@ 4 Binary Service
Under the binary service, there is no distortion in the service provided to consumers in the optimal mechanism. When the service is binary, the service mismatch if any is fixed at the horizontal differentiation level $\delta^{2}$. Distorting a service provision requires a large compensation to consumers. Instead, without a mismatch in service provision, the intermediary can extract the most consumer surplus by charging a higher service fee.
The obedience constraint in (8) for charging a high product price, the producer’s revenue without any data in (10), and the type-H consumers’ participation constraint in (13) depend on the market prior $\mu_{0}$. The following discusses by two cases: $\mu_{0}>v_{L}/v_{H}$ and $\mu_{0}<v_{L}/v_{H}$. [^17] The special parameter case of $\mu_{0}=v_{L}/v_{H}$ is captured by either case.
4.1 Intermediary Optimality for $\mu_{0}>v_{L}/v_{H}$
When $\mu_{0}>\frac{v_{L}}{v_{H}}$, the obedience constraint for the producer to charge a high product price $OB_{H}$ (8) is redundant: whenever (9) holds, (8) is also true. To see this, recall $\frac{v_{L}}{v_{H}}-\mu_{0}<0$ and it makes (8) redundant given (9). In other words, if the producer follows the intermediary’s low price recommendation, then the producer would also follow the high price recommendation. We only need to consider $OB_{L}$ (9) for $\mu_{0}>\frac{v_{L}}{v_{H}}$. In particular, if (9) is binding, then $\pi\left(l|H\right)=C\cdot\pi\left(l|L\right)$. [^18] As $\mu_{0}>v_{L}/v_{H}$, $C\in\left(0,1\right)$ from Notation 1.
@@ 4.1 Intermediary Optimality for $\mu_{0}>v_{L}/v_{H}$
The type-H consumers will be still charged a product price $p_{0}=v_{H}$ even if they reject the service offer, for $\mu_{0}>\frac{v_{L}}{v_{H}}$. Hence, the type-H consumers’ outside option $\max\left\{v_{H}-p_{0},0\right\}$ is zero. The participation constraint of the type-H consumers (13) is
@@
When $\mu_{0}>\frac{v_{L}}{v_{H}}$, the producer charges $v_{H}$ without buying any data. The type-H consumers purchase the product and the type-L consumers are excluded from the market. The producer’s revenue is $R_{0}=\mu_{0}v_{H}$. The producer’s net gain from data trade in (11) becomes
@@
The market inclusion effect of data trade is positive $(1-\mu_{0})\pi(l|L)>0$: The data trade extends the product market by allowing the producer to serve the type-L consumers (of mass $1-\mu_{0}$) with a higher probability $\pi(l|L)>0$. The surplus extraction effect of data trade is negative $-\mu_{0}\Delta v\pi(l|H)<0$. The producer extracts less type-H consumer surplus with data trade. With data trade, the producer charges the type-H consumers a low price $v_{L}$ with probability $\pi\left(l|H\right)$. Instead, without any data trade, all the type-H consumers are charged the high price $v_{H}$ and the intermediary does not yield any surplus to consumers.
The first main result on the intermediary’s optimal mechanism for data collecting and selling is summarized in the following proposition.
Proposition 1.
Suppose $\mu_{0}>v_{L}/v_{H}$. In the optimal mechanism, the intermediary sets the service fee to extract all the consumers surplus in both service and product markets
$f_{L}=V,\text{ }f_{H}=V+\pi\left(l|H\right)\Delta v$
The optimal data trade depends on the cross-market comparison between the mismatch cost in service $\delta^{2}$ and the vertical differentiation in the product $\Delta v$:
  1. 1.
    If $\delta^{2}\geq\Delta v$, the intermediary’s data trade is
    and the data trade fee is
    $T=-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)v_{L}.$
    In this case, social welfare is
    $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}v_{H}.$
  2. 2.
    If $\delta^{2}<\Delta v$, the intermediary’s data trade is
    $\pi\left(l|L\right)=\frac{\delta^{2}}{\Delta v},\text{ }\pi\left(l|H\right)\in\Big{[}0,\frac{C\delta^{2}}{\Delta v}\Big{]}$
    and the data trade fee is
    $T=-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}.$
    In this case, social welfare is
    $V+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}+\mu_{0}v_{H}.$
@@
The intermediary extracts all the surplus from both types of consumers in the service and product markets, the participation constraints of both types of consumers are binding. [^19] This is different from the standard mechanism design, where the type-H consumers have some surplus as information rent. With data trade, all consumers surplus are endogenous and extracted in the optimal mechanism. That is,
$f_{H}$ $=$ $V+\pi\left(l|H\right)\Delta v$
$f_{L}$ $=$ $V$
@@
where the profit from the service market is $V+\mu_{0}\pi\left(l|H\right)\Delta v$ and from the data market is $-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$. The surplus-shifting and the surplus-extracting offset each other. [^20] This is true because the intermediary has full bargaining power over the data trade surplus as it is a monopolist in the data market.
@@
With binding $IR_{L}$ (15) and $IR_{H}$ (17), then $IC_{L}$ (14) will be redundant, the intermediary only needs to consider $OB_{L}$ (9) and $IC_{H}$ (12). The intermediary’s reduced problem is to choose $\pi$ to solve
@@
In the first constraint: $\pi\left(l|L\right)\Delta v-\delta^{2}\leq 0$, if a type-H consumer reports the truth, she will get nothing. Instead, if she misreports, she will get the surplus $\Delta v$ with probability $\pi\left(l|L\right)$ but incur mismatch loss in the service $\delta^{2}$. To incentivize the type-Hs to tell the truth, the intermediary’s choice of $\pi\left(l|L\right)$ has to satisfy $\pi\left(l|L\right)\Delta v\leq\delta^{2}$.
@@
This reduced problem clarifies the tradeoff faced by the intermediary. The type-H consumers’ expected surplus in the product by misreporting, $\pi\left(l|L\right)\Delta v$ is bounded by the mismatch in the service $\delta^{2}$. However, the intermediary prefers $\pi\left(l|L\right)$ as high as possible because it can extract the type-L consumers’ surplus with a higher probability in the product market.
@@
For the intermediary, there are two instruments to incentivize the type-H consumers to tell the truth: decrease the probability of charging $v_{L}$ for those who report $L$, and the mismatch (wrong service) itself as a penalty. The mismatch is preferred by the intermediary because it does not require to lower $\pi\left(l|L\right)$. Hence, when the mismatch cost in service is high enough as a penalty, the type-H consumers will not misreport and the intermediary would like to set $\pi\left(l|L\right)$ as high as possible, that is $\pi\left(l|L\right)=1$. The type-L consumers are all served in the product market. In this case, the total profit of the intermediary is $V+\left(1-\mu_{0}\right)v_{L}$. Social welfare is the sum of consumers surplus 00, producer surplus $\mu_{0}v_{H}$ and intermediary surplus $V+\left(1-\mu_{0}\right)v_{L}$, which is $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}v_{H}$.

### v3

@@ 2 The Model
The intermediary plays in both service and data markets by offering a contract of data collection and data trade. In the service market the intermediary is a monopolist who provides consumers with services to collect their taste data. The intermediary sells data via a data offer and the producer buys data for product pricing decisions. The intermediary is the unique data source. Its profit comes from service provision and data sale.
By the revelation principle (Myerson, 1979), the intermediary designs a direct mechanism which specifies: (1) a service menu for consumers including the service provision $x:\{L,H\}\to\mathcal{X}$ where $\mathcal{X}$ is the set of service; and the service fee $f:\{L,H\}\to\mathbb{R}$; and (2) a data offer for the producer consisting of the disclosure policy captured by a pair of distributions over disclosed signals about consumers types $\pi:\left\{L,H\right\}\to\Delta\mathcal{S}$ where $\mathcal{S}$ is the set of possible signals; and the lump-sum data fee $T\in\mathbb{R}$. Consumers observe all the specifications. If a consumer reports $\theta^{\prime}$, then a signal will realize according to $\pi\left(s|\theta^{\prime}\right)$ in the product market. In sum, the intermediary offers a contract of data collection and data trade of which $\left(\mathcal{X},x,f\right)$ and $\left(\left\{\mathcal{S},\pi\right\},T\right)$ specify the part of service market and data market respectively.
@@
Consumers privately observe type $\theta$. The intermediary posts the entire contract $\{(\mathcal{X},x,f),(\{\mathcal{S},\pi\},T)\}$ including the service menu in the service market and data offer in the data market. Then consumers and the producer make acceptance decisions. If consumers accept, they report type $\theta^{\prime}$ and then get service $x_{\theta^{\prime}}$ and pay $f_{\theta^{\prime}}$. If consumers reject, they get the outside option in the service market. If the producer accepts, after data trade, in the product market a signal $s$ about consumer’s type is realized according to $\pi$. The producer observes $s$ and then sets the product price. Consumers decide whether to purchase the product or not. If the producer rejects the data offer, then the product pricing is made under the prior information $\mu_{0}$. The timing of events is summarized in Figure 2.
@@
A key modeling assumption is the perfect correlation between consumers’ taste and their willingness to pay. That is to say, $v_{L}\neq v_{H}$ in the setup. Otherwise, the data about consumers’ tastes is not valuable to the producer. The relation between consumers’ taste parameter and their willingness to pay can be represented by $\rho_{\theta}=\text{Pr}\left(v_{H}|\theta\right)=1-\text{Pr}\left(v_{L}|\theta\right)$. The assumption reads $(\rho_{H}=1$, $\rho_{L}=0)$ or $(\rho_{H}=0$, $\rho_{L}=1)$. If we relax the assumption by $\rho_{H}$, $\rho_{L}\in\left(0,1\right)$ with $\rho_{H}\neq\rho_{L}$, the privacy results don’t change. [^14] We describe in the supplementary online appendix how our results changed/remained unchanged as we move toward the imperfect correlation to consider heterogeneity in valuations conditional on type. Discussions on other assumptions are relegated to Section 6.
@@ 3 Preliminaries
In the product market, a consumer of $\theta$ will purchase the product if and only if $v_{\theta}\geq p$ at the product price $p$. After observing a signal $s\in\mathcal{S}$ from the traded data $\pi$, the producer’s posterior belief is $\mu_{s}\equiv\text{Pr}(\theta=H|s)$ or $\text{Pr}(v=v_{H}|s)$
@@
At the time of data trade, the producer’s expected product revenue from accepting the data trade offer $(\{\mathcal{S},\pi\},T)$ is
$R(\pi)=\mathbb{E}_{\pi}\left[\max\left\{v_{H}\mu_{s},v_{L}\right\}\right]$ (4)
The producer’s payoff is $R(\pi)-T$ if accepting the data offer.
@@
$u(\theta,\theta^{\prime};\pi)=V-\left(x_{\theta^{\prime}}-\theta\right)^{2}-f_{\theta^{\prime}}+\mathbb{E}_{\pi\left(s|\theta^{\prime}\right)}\left[\max\left\{v_{\theta}-p_{s},0\right\}\right]$ (5)
@@ 3.1 Producer Optimality
Plugging (2) into (6) (7), the obedience constraints for the producer to follow pricing recommendations are summarized into
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)\geq\max\left\{\frac{v_{L}}{v_{H}}-\mu_{0},0\right\}$ (8)
Now we turn to the producer’s acceptance decision over the intermediary’s data trade offer $\left(\left\{\mathcal{S},\pi\right\},T\right)$. When the producer does not accept the intermediary’s data offer, he has no more information than the prior $\mu_{0}$ and then the revenue is $R_{0}=\max\{v_{H}\mu_{0},v_{L}\}$. The producer accepts the data offer if and only if the product revenue from accepting is weakly larger than the one from rejecting
@@
where $\mathbbm{1}$ denotes the indicator function. The data trade has two effects on the producer’s revenue. The first is the change in the probability of serving the type-L consumers. With data trade the producer sells to the type-L consumers (of mass $(1-\mu_{0})$) with probability $\pi(l|L)$. Instead, without data trade the type-L consumers are served with probability $\mathbbm{1}_{\mu_{0}<\frac{v_{L}}{v_{H}}}$. This is the extensive market inclusion effect of data trade.
The second effect is the change in the probability of yielding surplus to the type-H consumers. With data trade the producer charges the type-H consumers (of mass $\mu_{0}$) a low price $v_{L}$ with probability $\pi(l|H)$. Without data trade the type-H consumers are charged the low price $v_{L}$ with probability $\mathbbm{1}_{\mu_{0}<\frac{v_{L}}{v_{H}}}$. This is the intensive surplus extraction effect of data trade. [^15] It is direct that when $\mu_{0}<\frac{v_{L}}{v_{H}}$, the market inclusion effect is negative: The market serving the type-L consumers shrinks. The surplus extraction effect is positive: The producer extracts the type-H consumer surplus with a larger probability. Instead, when $\mu_{0}>\frac{v_{L}}{v_{H}}$, the market inclusion effect is positive: The market serving the type-L consumers expands. the surplus extraction effect is negative: the producer extracts the type-H consumers surplus with a smaller probability.
The triangle gray set in Figure 3 illustrates the feasible data trade gain, which depends on the prior. Take $\mu_{0}>v_{L}/v_{H}$ as an example. Without data trade, the revenue is $\mu_{0}v_{H}$ which is at point $A$. It can go up to point $B$ under a data trade which induces posteriors $\mu_{l}$ and $\mu_{h}$. If the traded data is of full information, the intermediary extracts the maximum data trade surplus given $\mu_{0}$, which is $\mu_{0}v_{H}+\left(1-\mu_{0}\right)v_{L}$.
@@
Figure 3: The triangle gray part is the feasible data trade gain from a data trade. Given $\mu_{0}>v_{L}/v_{H}$: the product revenue is $\mu_{0}v_{H}$ at point AA without data trade. It goes up to point $B$ under a partial information data which induces posteriors $\mu_{l}$ and $\mu_{h}$. If the data is of full information, the product revenue is $\mu_{0}v_{H}+\left(1-\mu_{0}\right)v_{L}$.
@@ 3.2 Consumers Optimality
If the type-H consumers reject the service offer, they will get nothing in the service market. In this case, the producer will treat consumers as coming from the population and make product pricing decisions based on the prior $\mu_{0}$. [^16] The equilibrium selection criterion restricts off-path beliefs to the prior, which is in the spirit of “no signaling what you don’t know” (Fudenberg and Tirole, 1991): the intermediary cannot sell data to the producer if the offer is rejected by consumers, and the producer doesn’t update beliefs about consumers from a non-participation decision. See more discussions in Section 6. That is, the product price $p_{0}$ without extra data is determined by (3) with the belief $\mu_{0}$. So the type-H consumers’ continuation value of rejecting the intermediary’s contract is
@@ 3.3 Intermediary Optimality
The intermediary maximizes the total profit (1) from both the service and data markets by specifying a contract subject to the producer’s obedience constraints (8) and participation constraint (9), the consumers’ incentive constraints (11) (13) and participation constraints (12) (14), and the following two feasibility conditions of data trade
@@
4 Optimal Design
It is ready to characterize the intermediary optimal mechanism. Section 4.1 assumes the binary service to focus on the role of data trade; Section 4.2 returns to the general service to complete the characterization of intermediary optimal design of data collection and trade.
4.1 Special Case of Binary Service
To focus on the role of data trade, this subsection assumes a binary service $\mathcal{X}=\left\{L,H\right\}$, [^17] One can treat the assumption as a delegated provision of service by the intermediary to its department of marketing, while the data trade is directly controlled by the intermediary. The marketing department’s choice of service level does not coordinate with the data trade policy. This is one of the practices: the intermediary is usually not the producer of service, but recommends service. under which there is no distortion in the service provided to consumers in the intermediary optimal mechanism.
Lemma 2.
If $\mathcal{X}=\left\{L,H\right\}$, then $x_{H}=H$ and $x_{L}=L$.
We make the following short notations for exposures of results.
@@ 4.1 Special Case of Binary Service
The parameter $\delta^{2}$ captures the variation in consumers’ tastes, reflecting the horizontal differentiation level in the service market, while $\Delta v$ captures the variation in consumers’ willingness to pay, measuring the vertical differentiation level in the product market. When the service is binary, the service distortion if any is fixed at $\delta^{2}$. Distorting a service provision requires compensation to consumers. Instead, without any distortion of service provisions, the intermediary can extract the most consumer surplus by charging a higher service fee. Hence, the intermediary will not distort the service provided to consumers.
The intermediary’s optimal mechanism for data collecting and selling is summarized in the following proposition.
Proposition 1 (Optimal data—binary service).
Given a binary service space. The optimal data trade depends on the cross-market comparison between the horizontal differentiation level in the service $\delta^{2}$ and the vertical differentiation in the product $\Delta v$:
  1. 1.
    Suppose $\mu_{0}>v_{L}/v_{H}$. The intermediary’s data trade is $\pi\left(l|L\right)=\min\{\frac{\delta^{2}}{\Delta v},1\}$, $\pi\left(l|H\right)$ can be anywhere between 0 and $C\min\{\frac{\delta^{2}}{\Delta v},1\}$.
  2. 2.
    Suppose $\mu_{0}<v_{L}/v_{H}$. The intermediary’s optimal data trade is $\pi\left(l|L\right)=1$, $\pi\left(l|H\right)$ can be anywhere between $\max\left\{1-\frac{\delta^{2}}{\Delta v},0\right\}$ and $1$.
The obedience constraints of the producer in (8), the producer’s revenue without any data in (9), and the type-H consumers’ participation constraint in (12) depend on the market prior $\mu_{0}$. The following illustrates the results, taking $\mu_{0}>v_{L}/v_{H}$ as an example. [^18] The detailed results of $\mu_{0}<v_{L}/v_{H}$ are relegated to the appendix. The special parameter of $\mu_{0}=v_{L}/v_{H}$ is captured by either case, but it also raises more potential solutions. To see this, note that at $\mu_{0}=v_{L}/v_{H}$, a) the constraint (8) falls into the case of $\mu_{0}>v_{L}/v_{H}$; b) $p_{0}$ equals to $v_{L}$ or $v_{H}$ and $R_{0}=v_{L}$ or $v_{H}\mu_{0}$, but (9) doesn’t affect analysis because $T=R(\pi)-R_{0}$ and $R_{0}$ enters the objective as a constant; c) (12) becomes either (17) of the case of $\mu_{0}>v_{L}/v_{H}$, or (19) of the case of $\mu_{0}<v_{L}/v_{H}$. Thus, the special case of $\mu_{0}=v_{L}/v_{H}$ could raise more solutions as there is an extra combination of constraints including (16) and (19) together. This combination of constraints affects the data fee $f_{H}$, but does not affect the results about data trade. When $\mu_{0}>v_{L}/v_{H}$, the obedience constraint for the producer to charge a high price (6) is redundant and the obedience constraint for the producer to charge a low price reads
$\pi\left(l|H\right)\leq C\pi\left(l|L\right)$ (16)
@@
where $C\in\left(0,1\right)$ is in 1. The type-H consumers will be charged a product price $p_{0}=v_{H}$ when they reject the service offer, for $\mu_{0}>v_{L}/v_{H}$. Hence, the type-H consumers’ outside option $\max\left\{v_{H}-p_{0},0\right\}$ is zero. The participation constraint of the type-H consumers (12) is
@@
The producer charges $v_{H}$ without buying any data. The type-H consumers purchase the product and the type-L consumers are excluded from the market. The producer’s revenue is $R_{0}=\mu_{0}v_{H}$. The producer’s net gain from data trade in (10) becomes
@@
The market inclusion effect of data trade is positive $(1-\mu_{0})v_{L}>0$: The data trade extends the product market by allowing the producer to serve the type-L consumers (of mass $1-\mu_{0}$) with a higher probability $\pi(l|L)>0$. The surplus extraction effect of data trade is negative $\mu_{0}\Delta v<0$. The producer extracts less type-H consumer surplus with data trade. With data trade, the producer charges the type-H consumers a low price $v_{L}$ with probability $\pi\left(l|H\right)$. Instead, without any data trade, all the type-H consumers are charged the high price $v_{H}$ and the intermediary does not yield any surplus to consumers.
@@
The intermediary extracts all the surplus from both types of consumers in the service and product markets, the participation constraints of both types of consumers are binding. [^19] This is different from the standard mechanism design: all consumer surpluses are endogenous and extracted in the optimal mechanism with data trade. That is,
$f_{H}=V+\pi\left(l|H\right)\Delta v,\text{ }f_{L}=V.$
@@
where the profit from the service market is $V+\mu_{0}\pi\left(l|H\right)\Delta v$ and from the data market is $\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$. The surplus-shifting and the surplus-extracting offset each other. [^20] It is because the intermediary has full bargaining power over the data trade surplus in the data market. Section 5 instead looks at the partial bargaining power of the intermediary.
@@
With binding $IR_{L}$ (14) and $IR_{H}$ (17), then $IC_{L}$ (13) will be redundant, the intermediary only needs to consider $OB_{L}$ (16) and $IC_{H}$ (11). The intermediary’s reduced problem is to solve
@@
In the first constraint of $\pi\left(l|L\right)\Delta v-\delta^{2}\leq 0$, if a type-H consumer reports the truth, she will get nothing. Instead, if she misreports, she will get the surplus $\Delta v$ with probability $\pi\left(l|L\right)$ but incur mismatch loss in the service $\delta^{2}$. To incentivize the type-Hs to tell the truth, the intermediary’s choice of $\pi\left(l|L\right)$ has to satisfy $\pi\left(l|L\right)\Delta v\leq\delta^{2}$.
@@
The above reduced problem clarifies the tradeoff faced by the intermediary. The type-H consumers’ expected surplus in the product market by misreporting, $\pi\left(l|L\right)\Delta v$, is bounded by the mismatch in the service $\delta^{2}$. However, the intermediary prefers $\pi\left(l|L\right)$ as high as possible because it can extract the type-L consumers’ surplus with a higher probability in the product market.
@@
Intuitively, there are two instruments for the intermediary to incentivize the type-H consumers to tell the truth: lowering the probability of charging $v_{L}$ for those who report $L$, and the mismatch (wrong service) itself as a penalty. The mismatch penalty is preferred by the intermediary because it does not require to lower $\pi\left(l|L\right)$.
When the horizontal differentiation level is weakly above the vertical one, $\delta^{2}\geq\Delta v$, the mismatch cost in service is high enough as a penalty. In this case, the type-H consumers will not misreport and the intermediary would like to set $\pi\left(l|L\right)$ as high as possible, that is $\pi\left(l|L\right)=1$. The type-L consumers are all served in the product market.